OpenStudy (anonymous):

Find the percentile rank of 90

OpenStudy (anonymous):

Which of the following inequalities matches the graph? 1) 3x - 2y alt= 4 2) 3x - 4y alt= 2 3) 3x - 2y alt= 4 4) The correct inequality is not listed